log(x-3) + log (x) = 1
Answer by nerdybill(5404)

log(x-3) + log (x) = 1
log[(x-3)(x)] = 1
(x-3)(x) = 10^1
x^2-3x = 10
x^2-3x-10 = 0
(x-5)(x+2) = 0
.
x = {-2, 5}